OpenStudy (anonymous):

The plane (x/2)+(y/4)+(z/3)=1 intersects the x,y,and z axes in points P, Q, and R. FInd the Area of the triangle PQR @ganeshie8

ganeshie8 (ganeshie8):

familiar with cross product ?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

then it is easy :) you just need to know that `x/a + y/b + z/c = 1` is the intercept form of plane

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

ok... hold on. im trying to grasp it in my brain :|

ganeshie8 (ganeshie8):

Area of triangle PQR can be given by below expression \[\large \dfrac{\|PQ \times PR\|}{2}\]

OpenStudy (anonymous):

ok, again, so it doesnt matter what letter ( p,q,r ) you assign to x,y, or z?

ganeshie8 (ganeshie8):

you're right, it doesnt matter and it should make sense why it doesnt matter

ganeshie8 (ganeshie8):

remember the formula for area of triangle with base b and height h ? \[A = \frac{b*h}{2}\] you can pick ANY of the three sides for base and the corresponding altitude for height. the area doesn't depend on which side you pick for base